# AdaptGCD: Multi-Expert Adapter Tuning for Generalized Category Discovery

### v1

@@ 3.3 Multi-Expert Adapter
Basic Adapter. The designation of the adapter in this work follows the basic architecture proposed by AdaptFormer DBLP:conf/nips/adaptformerChenGTWSWL22 , which replaces the Feed-Forward Network (FFN) in the Transformer with the AdaptMLP module. This module contains two parallel sub-branches, one is identical to the Feed-Forward block while the other is an learnable lightweight bottleneck module. The bottleneck module includes a down-projection layer with parameters $\textbf{W}_{\rm down}\in\mathbb{R}^{\hat{d}\times d}$, $\textbf{b}_{\rm down}\in\mathbb{R}^{1\times\hat{d}}$ and an up-projection layer $\textbf{W}_{\rm up}\in\mathbb{R}^{d\times\hat{d}}$, $\textbf{b}_{\rm up}\in\mathbb{R}^{1\times{d}}$, where $d$ and $\hat{d}$ denote the input dimension of FFN and middle dimension of bottleneck ($\hat{d}<d$). The down-projection layer and the up-projection layer are connected with the ReLU layer for non-linear properties and the entire bottleneck module is linked to the original FFN with the residual connection, as depicted in Figure 1. Specifically, for the features of the $i$-th sample fed into FFN of the $l$-th Transformer blocks $\widetilde{\textbf{x}}_{l,i}$, the forward process of bottleneck module to yield the adapted feature $\Delta{\textbf{x}}_{l,i}$ is formulated as follows:
@@
Multiple-Expert Adapter. As previously mentioned DBLP:conf/iclr/CaoBL22 , data from older classes yields more supervisory information. It would result in the preference for the old class as the network receives more gradients from the old-class data than the new-class data. To counteract this implicit bias, the strategy of multi-expert technology is adopted to integrate with the adapter. This approach assigns different types of data to different expert adapters, therefore achieving separation for the data at the network level and alleviating the interference between new classes and old classes.
Multiple-expert adapter (MEA) consists of $T$ experts for handling different features and one route function to assign weights to these experts. More specifically, for the input feature $\widetilde{\textbf{x}}_{l,i}$, MEA leverages the following route function to estimate the weight $\boldsymbol{\omega}_{l,i}$ of multiple experts:
@@
where $\textbf{W}_{\rm route}\in\mathbb{R}^{T\times d}$ denotes the trainable parameters in the route function and $\tau_{r}$ indicates the tempertures. The forward process of the bottleneck module is rectified as Eq.(4):
@@
where $\textbf{W}_{\rm down}^{t},\textbf{b}_{\rm down}^{t},\textbf{W}_{\rm up}^{t}$ and $\textbf{b}_{\rm up}^{t}$ denotes the parameters of the $t$-expert. By this, the multi-expert mechanism enables the network to develop powerful capabilities to flexibly handle the data from new and old classes. According to previous research DBLP:journals/corr/abs-2402-08562 , allocating more adapter experts in higher blocks enhances the effectiveness of models compared with inserting them in the lower blocks. Hence, the MEA structure is incorporated only in the last $P$ blocks out of the total $L$ blocks.
3.4 Route Assignment Constraint
@@
In the multi-expert adapter, the route assignment constraint is required to supervise and control the route distribution. The assignment mainly focuses on two aspects: First, for all data, the load of all experts needs to be balanced to make full use of the resources of experts. Second, for data in old or new classes, the constraint assigns the corresponding experts to them so that the data can be well separated at the routing level. These two aspects correspond to the balanced load loss and the partial balanced load loss, which are introduced in this part.
@@
Probability for Route Assignment. It is worthy noting that, in order to obtain the sample-wise route assignment probability, it is necessary to combine the gate vectors of different tokens for the same sample. Assume that $\boldsymbol{\omega}_{l,i,v}$ denotes the gate probability from the $v$-th token of the $i$-th sample, within the $l$-th blocks, then the route assignment probability for the $i$-th sample ${\boldsymbol{\omega}}_{l,i}$ can be formulated as ${\boldsymbol{\omega}}_{l,i}={\rm Softmax}\left(\frac{1}{V}\sum_{v=1}^{V}\boldsymbol{\omega}_{l,i,v}/\tau_{g}\right)$, where $V$ denotes the number of tokens and $\tau_{g}$ is the temperature, and ${\rm Softmax}(\cdot)$ is utilized to sharpen the gate vectors.
@@
Balanced Load Loss. The balanced load loss is designed to ensure the maximal usage of diverse experts. Previous research on mixture-of-expert DBLP:journals/corr/loramoeabs-2312-09979 demonstrated a tendency towards a “winner-takes-all” DBLP:journals/corr/loramoeabs-2312-09979 situation, which indicates that few experts are frequently activated and domain the learning process, resulting in the waste of some experts. To mitigate this, in this study, the mean distribution of experts in the $l$-th layer ${\boldsymbol{\omega}}_{l}$, averaging the route assignment probabilities across all samples, is aligned to a uniform distribution $I$, where $I(t)=1/T,t=1,...T$. The formulation is depicted in Eq. (5):
${\boldsymbol{\omega}}_{l}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}{\boldsymbol{\omega}}_{l,i},\quad\mathcal{L}_{bl}=\sum_{l=L-P+1}^{L}\mathbb{D}_{\rm KL}\left({\boldsymbol{\omega}}_{l}||{I}\right),$ (5)
@@
Partial Balanced Load Loss. Considering the imbalance of supervision between new classes and old classes, we propose a partial balanced load loss to separate the new-class and old-class data into different experts and reduce their interference. Since the exact class label of the samples is not determined during training, the pseudo labels are utilized as a substitute. As shown in Eq. (6), the pseudo-labels of the samples are obtained according to the predictions.
@@
Then the route assignment probabilities of the experts for the old and new classes are estimated as ${\boldsymbol{\omega}}^{old}_{l}$ and ${\boldsymbol{\omega}}^{new}_{l}$ in Eq. (7).
${\boldsymbol{\omega}}^{old}_{l}=\frac{\sum_{i}{\boldsymbol{\omega}}_{l,i}\mathbbm{1}(\hat{y}_{i}\in\mathcal{Y}_{l})}{\sum_{i}\mathbbm{1}(\hat{y}_{i}\in\mathcal{Y}_{l})},\quad{\boldsymbol{\omega}}^{new}_{l}=\frac{\sum_{i}{\boldsymbol{\omega}}_{l,i}\mathbbm{1}(\hat{y}_{i}\in\mathcal{Y}_{n})}{\sum_{i}\mathbbm{1}(\hat{y}_{i}\in\mathcal{Y}_{n})}.$ (7)
@@
$I^{old}(t)=\begin{cases}\frac{1}{|\mathcal{T}^{old}|}&t\in\mathcal{T}^{old}\\ 0&t\in\mathcal{T}^{new}\\ \end{cases},I^{new}(t)=\begin{cases}0&t\in\mathcal{T}^{old}\\ \frac{1}{|\mathcal{T}^{new}|}&t\in\mathcal{T}^{new}\\ \end{cases}.$ (8)
These two distribution functions respectively describe the target assignment distribution of data in the old and new data. Then, as shown in Eq. (9), the Kullback-Leibler divergence is adopted to align ${\boldsymbol{\omega}}^{old}_{l}$ and ${\boldsymbol{\omega}}^{new}_{l}$ with the predefined target $I^{old}$ and $I^{new}$:
$\mathcal{L}_{pbl}=\sum_{l=L-1+P}^{L}\left(\mathbb{D}_{\rm KL}\left({\boldsymbol{\omega}}^{old}_{l}||{I}^{old}\right)+\mathbb{D}_{\rm KL}\left({\boldsymbol{\omega}}^{new}_{l}||{I}^{new}\right)\right).$ (9)
@@
In the final step, the route assignment loss $\mathcal{L}_{ra}$ for the AdaptGCD is collected as the weighted sum of the two losses, i.e. $\mathcal{L}_{ra}=\beta\mathcal{L}_{bc}+\alpha\mathcal{L}_{pbl}$, where $\alpha,\beta$ are the balancing factors.
@@
Overall Loss. The overall loss of the AdaptGCD is collected as the sum of SimGCD loss and the route assignment loss, as formulated in Eq. (10):
$\mathcal{L}_{overall}=\mathcal{L}_{gcd}+\mathcal{L}_{ra}.$ (10)
4 Experiments
4.1 Experimental Setup
@@
Datasets. The effectiveness of the proposed method is validated on three characteristic benchmark: the generic image recognition benchmark including CIFAR10/100 krizhevsky2009learning and ImageNet-100 DBLP:conf/eccv/TianKI20 , the semantic shift benchmark including CUB-200 welinder2010caltech , Stanford Cars krause20133d , and FGVC-Aircraft DBLP:journals/corr/MajiRKBV13 and the harder long-tail dataset Herbarium 19 tan2019herbarium . Following SimGCD DBLP:conf/iccv/simgcdWenZQ23 , a subset of all the classes $\mathcal{Y}^{l}$ is sampled as the old classes while 50% of the images from the old classes are selected to construct the labeled set $\mathcal{D}^{l}$. The rest of the images are collected as the unlabeled set $\mathcal{D}^{u}$.
@@
Implementation details and evaluations. Following GCD DBLP:conf/cvpr/gcdVazeHVZ22 , the proposed method is trained with a ViT-B/16 backbone pre-trained with DINO. The output of [CLS] tokens is utilized as the feature of the image. For the training details, the AdaptGCD is trained with a batch size of 128 for 200 epochs with an initial learning rate of 0.1 decayed with a cosine schedule on each dataset. For the adapter configuration, we utilize 8 experts with the bottleneck dimension $\hat{d}$ set as 64. The number of adapted blocks $P$ is assigned to be 6 on CUB-200, Scars and CIFAR10. As for the Aircraft, CIFAR100, Herbarium19 and Imagenet100, which present greater challenges for the classification, $P$ is designated to be 8. For expert assignments, 4 experts are utilized as old-class experts while another 4 experts are selected as new-class experts, i.e. $|\mathcal{T}^{old}|=|\mathcal{T}^{new}|=4$. The performance is evaluated with the clustering accuracy for old, new and all classes as DBLP:conf/cvpr/gcdVazeHVZ22 . It is also noted that we conduct the experiments in Table 1 and 2 for 3 independent runs and report the average.
@@
Table 1: Evaluation results on the generic image recognition benchmark. Bold values represent the best results, while underlined values denote the second best results.
Methods CIFAR10 CIFAR100 ImageNet100
@@
$k\mbox{-}means$ DBLP:conf/soda/ArthurV07 83.6 85.7 82.5 52.0 52.2 50.8 72.7 75.5 71.3
RankStats+ DBLP:journals/pami/HanREVZ22 46.8 19.2 60.5 58.2 77.6 19.3 37.1 61.6 24.8
UNO+ DBLP:conf/iccv/FiniSLZN021 68.6 98.3 53.8 69.5 80.6 47.2 70.3 95.0 57.9
ORCA DBLP:conf/iclr/CaoBL22 81.8 86.2 79.6 69.0 77.4 52.0 73.5 92.6 63.9
GCD DBLP:conf/cvpr/gcdVazeHVZ22 91.5 97.9 88.2 73.0 76.2 66.5 74.1 89.8 66.3
DCCL DBLP:conf/cvpr/PuZS23 96.3 96.5 96.9 75.3 76.8 70.2 80.5 90.5 76.2
SimGCD DBLP:conf/iccv/simgcdWenZQ23 97.1 95.1 98.1 80.1 81.2 77.8 83.0 93.1 77.9
PromptCAL DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 97.9 96.6 98.5 81.2 84.2 75.3 83.1 92.7 78.3
SPTNet wang2024sptnet 97.3 95.0 98.6 81.3 84.3 75.6 85.4 93.2 81.4
AdaptGCD 97.9 95.5 99.0 84.0 86.2 79.7 86.4 94.9 82.1
Table 2: Evaluated results on the semantic shift benchmark and Herbarium19. Bold values represent the best results, while underlined values represent the second best results.
Methods CUB-200 Scars Aircraft Herbarium19
@@
$k\mbox{-}means$ DBLP:conf/soda/ArthurV07 34.3 38.9 32.1 12.8 10.6 13.8 16.0 14.4 16.8 13.0 12.2 13.4
RankStats+ DBLP:journals/pami/HanREVZ22 33.3 51.6 24.2 28.3 61.8 12.1 26.9 36.4 22.2 27.9 55.8 12.8
UNO+ DBLP:conf/iccv/FiniSLZN021 35.1 49.0 28.1 35.5 70.5 18.6 40.3 56.4 32.2 28.3 53.7 14.7
ORCA DBLP:conf/iclr/CaoBL22 35.3 45.6 30.2 23.5 50.1 10.7 22.0 31.8 17.1 35.4 51.0 27.0
GCD DBLP:conf/cvpr/gcdVazeHVZ22 51.3 56.6 48.7 39.0 57.6 29.9 45.0 41.1 46.9 20.9 30.9 15.5
DCCL DBLP:conf/cvpr/PuZS23 63.5 60.8 64.9 43.1 55.7 36.2 - - - - - -
SimGCD DBLP:conf/iccv/simgcdWenZQ23 60.3 65.6 57.7 53.8 71.9 45.0 54.2 59.1 51.8 44.0 58.0 36.4
PromptCAL DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 62.9 64.4 62.1 50.2 70.1 40.6 52.2 52.2 52.3 37.0 52.0 28.9
SPTNet wang2024sptnet 65.8 68.8 65.1 59.0 79.2 49.3 59.3 61.8 58.1 43.4 58.7 35.2
AdaptGCD 68.8 74.5 65.9 62.7 80.6 54.0 57.9 65.2 54.3 45.7 60.6 37.7
4.2 Comparison with State-of-the-Arts

### v2

@@ IV-A Multi-Expert Adapter
1) Basic Adapter: The designation of the adapter in this work follows the basic architecture proposed by AdaptFormer [22], which replaces the Feed-Forward Network (FFN) in the Transformer with the AdaptMLP module. This module contains two parallel sub-branches, one is identical to the Feed-Forward block while the other is an learnable lightweight bottleneck module. The bottleneck module includes a down-projection layer with parameters $\textbf{W}_{\rm down}\in\mathbb{R}^{\hat{d}\times d}$, $\textbf{b}_{\rm down}\in\mathbb{R}^{1\times\hat{d}}$ and an up-projection layer $\textbf{W}_{\rm up}\in\mathbb{R}^{d\times\hat{d}}$, $\textbf{b}_{\rm up}\in\mathbb{R}^{1\times{d}}$, where $d$ and $\hat{d}$ denote the input dimension of FFN and middle dimension of bottleneck ($\hat{d}<d$). The down-projection layer and the up-projection layer are connected with the ReLU layer for nonlinear properties, and the entire bottleneck module is linked to the original FFN with the residual connection, as depicted in Fig. 2. Specifically, for the features of the $i$-th sample fed into FFN of the $l$-th Transformer blocks $\widetilde{\textbf{x}}_{l,i}$, the forward process of bottleneck module to yield the adapted feature $\Delta{\textbf{x}}_{l,i}$ is formulated as follows:
@@
2) Multiple-Expert Adapters: As previously mentioned [12], data from older classes yields more supervisory information. It would result in the preference for the old class as the network receives more gradients from the old-class data than the new-class data. To counteract this implicit bias, the strategy of multi-expert technology is adopted to integrate with the adapter. This approach assigns different types of data to different expert adapters, therefore achieving separation for the data at the network level and alleviating the interference between new classes and old classes.
Multiple-expert adapter (MEA) consists of $T$ experts for handling different features and a route function to assign weights to these experts. More specifically, for the input feature $\widetilde{\textbf{x}}_{l,i}$, MEA leverages the following route function to estimate the weight $\boldsymbol{\omega}_{l,i}$ of multiple experts:
@@
where $\textbf{W}_{\rm route}\in\mathbb{R}^{T\times d}$ denotes the trainable parameters in the route function and $\tau_{r}$ indicates the gating temperatures. The forward process of the bottleneck module is rectified as Eq.(7):
@@
where $\textbf{W}_{\rm down}^{t},\textbf{b}_{\rm down}^{t},\textbf{W}_{\rm up}^{t}$ and $\textbf{b}_{\rm up}^{t}$ denotes the parameters of the $t$-expert, respectively. More detailed implementation of MEA is in the Appendix. By this, the multi-expert mechanism enables the network to develop the powerful capabilities to flexibly handle the data from new and old classes. According to the previous research [61], allocating more adapter experts in higher blocks enhances the effectiveness of models compared with inserting them in the lower blocks. Hence, the MEA structure is incorporated only in the last $P$ blocks out of the total $L$ blocks.
IV-B Route Assignment Constraint
@@
In the multi-expert adapter, the route assignment constraint is required to supervise and control the route distribution. The assignment focuses mainly on two aspects: First, for all data, the assignment for all experts needs to be balanced to make full use of the resources of experts. Second, for data in old or new classes, the constraint assigns the corresponding experts to them so that the data can be well separated at the routing level. These two aspects correspond to the balanced assignment loss and the category-aware balanced assignment loss, which are introduced in this part. The proposal applies an assignment constraint on the average expert assignment of new and old class samples, rather than on individual samples as in the constraint from previous works [68]. This enables more flexible expert assignments and reduces sensitivity to inaccurate pseudo labels. More comparisons between the proposed route assignment constraint and the loss in previous works are depicted in the Appendix.
@@
1) Sample-Wise Probability for Route Assignment: It is worth noting that, in order to obtain the sample-wise route assignment probability, it is necessary to combine the gate vectors of different tokens for the same sample. Assume that $\boldsymbol{\omega}_{l,i,v}$ denotes the gate probability from the $v$-th token of the $i$-th sample within the $l$-th blocks, then the route assignment probability for the $i$-th sample ${\boldsymbol{\omega}}_{l,i}$ can be formulated as ${\boldsymbol{\omega}}_{l,i}={\rm Softmax}\left(\frac{1}{V}\sum_{v=1}^{V}\boldsymbol{\omega}_{l,i,v}/\tau_{g}\right)$, where $V$ denotes the number of tokens and $\tau_{g}$ is the routing temperature, and ${\rm Softmax}(\cdot)$ is utilized to sharpen the gate vectors.
@@
2) Balanced Assignment Loss: The balanced assignment loss is designed to ensure the maximal usage of diverse experts. Previous research on mixture-of-expert [68] demonstrated a tendency towards a “winner-takes-all” [68] situation, which indicates that few experts are frequently activated and domain the learning process, resulting in the waste of some experts. To mitigate this, in this study, the mean distribution of experts in the $l$-th layer ${\boldsymbol{\omega}}_{l}$, averaging the route assignment probabilities across all samples, is aligned to a uniform distribution $I$, where $I(t)=1/T,t=1,...T$. The formulation is depicted in Eq. (8):
${\boldsymbol{\omega}}_{l}=$ $\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}{\boldsymbol{\omega}}_{l,i},$ (8)
$\mathcal{L}_{ba}=$ $\sum_{l=L-P+1}^{L}\mathbb{D}_{\rm KL}\left({\boldsymbol{\omega}}_{l}||{I}\right),$
@@
3) Category-Aware Balanced Assignment Loss: Considering the imbalance of supervision between new classes and old classes, we propose a category-aware balanced assignment loss to separate new-class and old-class data into different experts and reduce their interference. Since the exact class label of the samples is not determined during training, the pseudo labels are utilized as a substitute. As shown in Eq. (9), pseudo labels of the samples are obtained according to the predictions.
@@
where $p_{i,k}$ indicates the predictive probability of $i$-sample in $k$-class. Then the route assignment probabilities of the experts for the old and new classes are estimated as ${\boldsymbol{\omega}}^{old}_{l}$ and ${\boldsymbol{\omega}}^{new}_{l}$ in Eq. (10).
${\boldsymbol{\omega}}^{old}_{l}=$ $\frac{\sum_{i}{\boldsymbol{\omega}}_{l,i}\mathbbm{1}(\hat{y}_{i}\in\mathcal{Y}_{l})}{\sum_{i}\mathbbm{1}(\hat{y}_{i}\in\mathcal{Y}_{l})},$ (10)
${\boldsymbol{\omega}}^{new}_{l}=$ $\frac{\sum_{i}{\boldsymbol{\omega}}_{l,i}\mathbbm{1}(\hat{y}_{i}\in\mathcal{Y}_{n})}{\sum_{i}\mathbbm{1}(\hat{y}_{i}\in\mathcal{Y}_{n})}.$
@@
$I^{old}(t)=$$\begin{cases}\frac{1}{|\mathcal{T}^{old}|}&t\in\mathcal{T}^{old}\\ 0&t\in\mathcal{T}^{new}\\ \end{cases},$ (11)
$I^{new}(t)=$$\begin{cases}0&t\in\mathcal{T}^{old}\\ \frac{1}{|\mathcal{T}^{new}|}&t\in\mathcal{T}^{new}\\ \end{cases}.$
These two distribution functions respectively describe the target assignment distribution of data in the old and new data. To enhance the separation of information between new and old classes, we employ the Kullback-Leibler divergence loss as a constraint, as shown in Eq. (12), to align the average probability of old and new classes ${\boldsymbol{\omega}}^{old}_{l}$ and ${\boldsymbol{\omega}}^{new}_{l}$ with the predefined targeted distribution $I^{old}$ and $I^{new}$. Unlike previous works [66] using weighted variance coefficients as a soft constraints to balance expert allocation within samples, this strategy emphasizes isolating average expert allocation of new class samples from old class samples through distribution alignment.
$\mathcal{L}_{cba}=\sum_{l=L-1+P}^{L}\left(\mathbb{D}_{\rm KL}\left({\boldsymbol{\omega}}^{old}_{l}||{I}^{old}\right)+\mathbb{D}_{\rm KL}\left({\boldsymbol{\omega}}^{new}_{l}||{I}^{new}\right)\right).$ (12)
@@
In the final step, the route assignment loss $\mathcal{L}_{ra}$ for the AdaptGCD is collected as the weighted sum of the two losses, i.e. $\mathcal{L}_{ra}=\beta\mathcal{L}_{ba}+\alpha\mathcal{L}_{cba}$, where $\alpha,\beta$ are the balancing factors.
@@
4) Overall Loss: The overall loss of AdaptGCD is collected as the sum of SimGCD loss and the route assignment loss, as formulated in Eq. (13):
$\mathcal{L}_{overall}=\mathcal{L}_{sgcd}+\mathcal{L}_{ra}.$ (13)
The proposal could also serve as a plug-and-play component for integration with other methods. In this study, we combine it with the advanced method SPTNet [14]. Specifically, SPTNet consists of two training stages that learn the parameters from the prompt and the backbone, respectively. To enhance its performance, we replace the second stage, which trains the backbone using partial tuning, with the proposed AdaptGCD training strategy.
V Experiments
V-A Experimental Setup
@@
1) Datasets: The effectiveness of the proposed method is validated on three characteristic benchmarks: the generic image recognition benchmark including CIFAR10/100 [69] and ImageNet-100 [70], the semantic shift benchmark including CUB-200 [71], Stanford Cars [72], and FGVC-Aircraft [73] and the harder long-tailed dataset Herbarium 19 [74]. Following SimGCD [55], a subset of all the classes $\mathcal{Y}^{l}$ is sampled as the old classes while 50% of the images from the old classes are selected to construct the labeled set $\mathcal{D}^{l}$. The rest of the images are collected as the unlabeled set $\mathcal{D}^{u}$. Furthermore, in CIFAR100, 80% of categories are assigned as old classes while the rest are considered as new ones. For the rest datasets, the portation of old classes is 50%. The basic settings of the datasets [13] are displayed in Table I.
@@
2) Implementation Details: Following GCD [13], the proposed method is trained with a ViT-B/16 backbone pre-trained with DINO. The output of [CLS] tokens is utilized as the feature of the image. For the training details, the AdaptGCD is trained with a batch size of 128 for 200 epochs with an initial learning rate of 0.1 decayed with a cosine schedule on each dataset. For the adapter configuration, we utilize 8 experts with the bottleneck dimension $\hat{d}$ set as 64.
TABLE I: Statistic information of the 7 datasets.
Benchmark Datasets Balance Labeled Unlabeled
# Image # Class # Image # Class
Semantic Shift Benchmark CUB-200 [71] 1.5K 100 4.5K 200
Aircraft [73] 1.7K 50 5.0K 100
Stanford Cars [72] 2.0K 98 6.1K 196
Generic Image Recognition Benchmark CIFAR10 [69] 12.5K 5 37.5K 10
CIFAR100 [69] 20.0K 80 30.0K 100
ImageNet100 [70] 31.9K 50 95.3K 100
Long-tailed Herbarium19 [74] 8.9K 341 25.4K 683
TABLE II: Detailed hyperparamter configuration of AdaptGCD on 7 datasets
Benchmark Datasets Configuration
$s$ $P$ $T$ $\hat{d}$ $\alpha$ $\beta$ $\tau_{r}$ $\tau_{g}$
Semantic Shift Benchmark CUB-200 [71] 0.4 6 8 64 0.03 0.1 5 0.1
Aircraft [73] 0.4 8 8 64 0.1 0.1 10 0.1
Scars [72] 0.4 6 8 64 0.1 0.1 10 0.1
Generic Image Recognition Benchmark CIFAR10 [69] 0.2 6 8 64 0.05 0.05 10 0.1
CIFAR100 [69] 0.8 8 8 64 0.05 0.05 10 0.1
ImageNet100 [70] 0.4 8 8 64 0.06 0.2 10 0.1
Long-tailed Herbarium19 [74] 0.4 8 8 64 0.05 0.05 10 0.1
The number of adapted blocks $P$ is assigned to be 6 on CUB-200, Scars and CIFAR10. As for the Aircraft, CIFAR100, Herbarium19 and Imagenet100, which present greater challenges for the classification, $P$ is designated to be 8. With these hyper-parameters, the proposed method is equipped with few trainable parameters in the backbone, totaling less than 6.4M, which is lower than the 7M in the last block trained in SimGCD, thus ensuring fairness. More analysis about the number of tunable parameters is presented on Appendix. For expert assignments, 4 experts are utilized as old-class experts while another 4 experts are selected as new-class experts, i.e. $|\mathcal{T}^{old}|=|\mathcal{T}^{new}|=4$. Further, Table II shows the rest hyper-parameter settings for the 7 datasets. For the integration of the proposed method and SPTNet, we follow the basic hyperparameter settings of both methods and conduct experiments on 5 datasets due to limited computational resources. More specific settings of hybrid methods are also detailed in the Appendix.
3) Evaluation: The performance is evaluated with the clustering accuracy for old, new and all classes as [13]. Given the ground truth $y_{i}$ and the predicted labels $\hat{y}_{i}$ for each $x_{i}$, the accuracies for the old/new/all classes $acc^{old}/acc^{new}/acc^{all}$ are formulated as Eq. (14).
$acc^{old}=$ $\frac{1}{|\mathcal{D}^{u,old}|}\sum_{i=1}^{|\mathcal{D}^{u,old}|}\mathbbm{1}(y_{i}={\rm Purmute}(\hat{y}_{i}))$ (14)
$acc^{new}=$ $\frac{1}{|\mathcal{D}^{u,new}|}\sum_{i=1}^{|\mathcal{D}^{u,new}|}\mathbbm{1}(y_{i}={\rm Purmute}(\hat{y}_{i}))$
$acc^{all}=$ $\frac{1}{|\mathcal{D}^{u}|}\sum_{i=1}^{|\mathcal{D}^{u}|}\mathbbm{1}(y_{i}={\rm Permute}(\hat{y}_{i}))$
where the $\mathcal{D}^{u,old}$ and $\mathcal{D}^{u,new}$ are the subsets of $\mathcal{D}^{u}$ only containing old and new classes. $\rm Permute(\cdot)$ is the optimal permutation matching the predicted cluster assignments to the ground truth class labels.
@@
TABLE III: Evaluation results on the generic image recognition benchmark. Bold values represent the best results, while underlined values denote the second best results.
Methods Tuning Strategy Generic Image Recognition Benchmark
CIFAR10 CIFAR100 ImageNet100
@@
$k\mbox{-}means$ [75] - 83.6 85.7 82.5 52.0 52.2 50.8 72.7 75.5 71.3
RankStats+ [76] Partial 46.8 19.2 60.5 58.2 77.6 19.3 37.1 61.6 24.8
UNO+ [77] Partial 68.6 98.3 53.8 69.5 80.6 47.2 70.3 95.0 57.9
ORCA [12] Partial 81.8 86.2 79.6 69.0 77.4 52.0 73.5 92.6 63.9
GCD [13] Partial 91.5 97.9 88.2 73.0 76.2 66.5 74.1 89.8 66.3
DCCL [53] Partial 96.3 96.5 96.9 75.3 76.8 70.2 80.5 90.5 76.2
SimGCD [55] Partial 97.1 95.1 98.1 80.1 81.2 77.8 83.0 93.1 77.9
CMS [78] Partial - - - 82.3 85.7 75.5 84.7 95.6 79.2
AMEND [79] Partial 96.8 94.6 97.8 81.0 79.9 83.3 83.2 92.9 78.3
PromptCAL [21] Prompt 97.9 96.6 98.5 81.2 84.2 75.3 83.1 92.7 78.3
GCA+PromptCAL [80] Prompt 95.5 95.9 95.2 82.4 85.6 75.9 82.8 94.1 77.1
SPTNet [14] Prompt 97.3 95.0 98.6 81.3 84.3 75.6 85.4 93.2 81.4
AdaptGCD Adapter 97.9 95.5 99.0 84.0 86.2 79.7 86.4 94.9 82.1
AdaptGCD+SPTNet Adapter 97.7 95.2 99.0 83.0 86.1 76.8 - - -
TABLE IV: Evaluation results on the semantic shift benchmark and long-tailed dataset. Bold values represent the best results, while underlined values denote the second best results.
Methods Tuning Strategy Semantic Shift Benchmark Long-tailed Dataset
CUB-200 Scars Aircraft Herbarium19
@@
$k\mbox{-}means$ [75] - 34.3 38.9 32.1 12.8 10.6 13.8 16.0 14.4 16.8 13.0 12.2 13.4
RankStats+ [76] Partial 33.3 51.6 24.2 28.3 61.8 12.1 26.9 36.4 22.2 27.9 55.8 12.8
UNO+ [77] Partial 35.1 49.0 28.1 35.5 70.5 18.6 40.3 56.4 32.2 28.3 53.7 14.7
ORCA [12] Partial 35.3 45.6 30.2 23.5 50.1 10.7 22.0 31.8 17.1 35.4 51.0 27.0
GCD [13] Partial 51.3 56.6 48.7 39.0 57.6 29.9 45.0 41.1 46.9 20.9 30.9 15.5
DCCL [53] Partial 63.5 60.8 64.9 43.1 55.7 36.2 - - - - - -
SimGCD [55] Partial 60.3 65.6 57.7 53.8 71.9 45.0 54.2 59.1 51.8 44.0 58.0 36.4
CMS [78] Partial 68.2 76.5 64.0 56.9 76.1 47.6 56.0 63.4 52.3 36.4 54.9 26.4
AMEND [79] Partial 64.9 75.6 59.6 56.4 73.3 48.2 52.8 61.8 48.3 44.2 60.5 35.4
$\mu$GCD [49] Partial 65.7 68.0 64.6 56.5 68.1 50.9 53.8 55.4 53.0 - - -
PromptCAL [21] Prompt 62.9 64.4 62.1 50.2 70.1 40.6 52.2 52.2 52.3 37.0 52.0 28.9
GCA+PromptCAL [80] Prompt 68.8 73.4 66.6 54.4 72.1 45.8 52.0 57.1 49.5 - - -
SPTNet [14] Prompt 65.8 68.8 65.1 59.0 79.2 49.3 59.3 61.8 58.1 43.4 58.7 35.2
AdaptGCD Adapter 68.8 74.5 65.9 62.7 80.6 54.0 57.9 65.2 54.3 45.7 60.6 37.7
AdaptGCD+SPTNet Adapter 71.6 75.7 69.6 63.5 75.7 57.6 60.8 68.1 57.2 - - -
V-B Comparison with State-of-the-Arts